A motel in Florida charges $90 per day for double occupancy and $80 for single occupancy. If 80 rooms are occupied for a total of $6930, how many rooms of each kind are occupied?.(please put the problem in system equation form which is {Ax+By=z
Gx+Hy=c
before you compute them.
thank you

Answer by algebrahouse.com(1659)

x = number of $80 per day rooms
y = number of $90 per day rooms

x + y = 80 {total number of occupied rooms is 80}
80x + 90y = 6930 {value of room times number of rooms equals total value}

-80x - 80y = -6400 {multiplied top equation by -80}
80x + 90y = 6930 {bottom original equation}
----------------
10y = 530 {added the two equations together}
y = 53 $90 per day rooms {divided each side by 10}

x + y = 80 {top original equation}
x + 53 = 80 {substituted 53, in for y, into top original equation}
x = 27 $80 per day rooms {subtracted 53 from each side}

53 $90 per day rooms
and
27 $80 per day rooms

For more help from me, visit: www.algebrahouse.com